# Remember, but also, Forget: Bridging Myopic and Perfect Recall Fairness with Past-Discounting

### v1

@@ Abstract.
Dynamic resource allocation in multi-agent settings often requires balancing efficiency with fairness over time—a challenge inadequately addressed by conventional, myopic fairness measures. Motivated by behavioral insights that human judgments of fairness evolve with temporal distance, we introduce a novel framework for temporal fairness that incorporates past-discounting mechanisms. By applying a tunable discount factor to historical utilities, our approach interpolates between instantaneous and perfect-recall fairness, thereby capturing both immediate outcomes and long-term equity considerations. Beyond aligning more closely with human perceptions of fairness, this past-discounting method ensures that the augmented state space remains bounded, significantly improving computational tractability in sequential decision-making settings. We detail the formulation of discounted-recall fairness in both additive and averaged utility contexts, illustrate its benefits through practical examples, and discuss its implications for designing balanced, scalable resource allocation strategies.
@@
Resource Allocation, Fairness, Multi-Agent RL
@@ 1. Introduction
Dynamic resource allocation is central to many real-world applications—from matching passengers to taxis Shah et al. (2020); Alonso-Mora et al. (2017) and distributing aid to the homeless Kube et al. (2019), to allocating life-saving vaccines Alamdari et al. (2024). In these settings, decision-makers must balance diverse agent preferences against resource constraints to maximize cumulative utility. Traditional fairness approaches impose constraints at each decision step (e.g., maximin fairness) but ignore the temporal dimension by treating each allocation in isolation. Such myopic methods fail to account for the evolving nature of fairness when past decisions and future opportunities are involved.
Recent work in multi-agent learning addresses temporal fairness by evaluating cumulative or terminal utilities Jiang and Lu (2019); Zimmer et al. (2021); Alamdari et al. (2024). Yet, these approaches either assume perfect recall—giving equal weight to all past allocations—or assess each allocation independently. In contrast, insights from behavioral economics and moral psychology reveal that human perceptions of fairness evolve over time. For example, events in the distant past tend to be perceived more abstractly Trope and Liberman (2010), and individuals naturally devalue outcomes that are further removed in time Frederick et al. (2002). Empirical studies also indicate that forgiveness increases as the temporal distance from a transgression grows Li et al. (2021); Wohl and McGrath (2007).
Motivated by these findings, we propose incorporating past discounting mechanisms into dynamic resource allocation. By discounting historical utilities, our approach offers a principled compromise between instantaneous and perfect-recall fairness. This method not only aligns more closely with observed human behavior but also ensures that the augmented state space remains bounded—a critical property for the convergence and scalability of reinforcement learning algorithms.
In this paper, we:
  • Highlight the limitations of traditional fairness approaches in dynamic settings.
  • Propose a framework that incorporates past discounting to balance short-term allocations with long-term fairness considerations.
  • Provide theoretical insights demonstrating how past discounting bounds the state space, thereby improving computational tractability.
By integrating theoretical analysis with behavioral insights, we aim to provide a strong argument for using past discounts for fair resource allocation in dynamic environments.
@@ 2. Related Work
A considerable body of research has explored fairness in resource allocation from both economic and algorithmic perspectives. Traditional fair division literature has focused on static fairness notions such as proportionality, envy‐freeness, and maximin share guarantees Budish (2011); Procaccia and Wang (2014). These approaches typically consider a one-shot allocation problem, enforcing fairness at each individual decision point. However, in dynamic or sequential settings—such as taxi matching, aid distribution, or vaccine allocation—the myopic application of static fairness criteria neglects the evolution of cumulative utilities over time.
Recent work in multi-agent reinforcement learning has begun to address fairness in dynamic settings. Several approaches enforce fairness by constraining per-step allocations or by evaluating the accumulated utilities at an intermediate or terminal stage Jiang and Lu (2019); Zimmer et al. (2021). More recently, Alamdari et al. (2024) introduce non‐Markovian fairness frameworks that explicitly incorporate past data by augmenting the state-space with historical allocations, looking at concepts like long‐term, anytime, and bounded fairness.
Beyond resource allocation, sequential fairness is also explored in multi-issue voting and online fair division. In these settings, fairness must account not only for a single decision but for a series of interdependent choices—ranging from perpetual voting schemes Lackner (2020) to online food bank allocation protocols Aleksandrov et al. (2015). Such cross-disciplinary work further highlights that static fairness notions must be adapted to dynamic environments if they are to reflect long-term, evolving perceptions of fairness.
While these models capture the temporal nature of decision-making, they typically assume either perfect recall (where past allocations are fully aggregated) or evaluate each allocation in isolation without considering how agents discount historical outcomes. In contrast, our work proposes a third paradigm—incorporating past-discounting mechanisms—motivated by evidence from behavioral economics and moral psychology.
Behavioral studies indicate that human fairness judgments are sensitive to temporal distance. Construal level theory posits that events in the distant past are represented more abstractly and evoke less emotional intensity Trope and Liberman (2010). Empirical research on time discounting shows that individuals systematically devalue outcomes as they recede into the past Frederick et al. (2002). These findings, along with studies on forgiveness in intergroup contexts Li et al. (2021); Wohl and McGrath (2007), imply that fairness assessments in real-world scenarios may benefit from discounting earlier allocations. This concept is reminiscent of the temporal discounting used in reinforcement learning Mnih et al. (2013); Sutton and Barto (1998), yet its explicit integration into fairness metrics for resource allocation remains largely underexplored.
Our work is also related to recent advances in fair multi-agent learning mechanisms. For example, methods for ride-hailing applications incorporate forecasting to balance future utility against current fairness Kumar et al. (2023), and hierarchical frameworks have been proposed to mediate between efficiency and fairness in multi-agent settings Jiang and Lu (2019). However, these approaches do not systematically account for the decay in the perceived value of past allocations. Instead, they either assume that all past outcomes are equally important (perfect recall) or that each allocation is evaluated in isolation (myopic fairness).
In summary, while the literature offers robust methods for achieving either instantaneous fairness or cumulative fairness with perfect recall, a practical middle ground remains unexplored. Our work fills this gap by introducing a past discounting framework that integrates historical context into fairness evaluation while accounting for the human tendency to weigh recent outcomes more heavily, all while keeping the state space computationally tractable.
@@ 3. Preliminaries
    $W_{N}(Z)=\prod_{i=1}^{n}z_{i},$ (4)
  • Generalized Gini Welfare, which is a family of functions that applies rank-based weights to the agent utilities, offering a flexible approach to balancing equity and efficiency Atkinson et al. (1970); Lorenz (1905); Zimmer et al. (2021).
@@
In many formulations, the allocation at time $t$, denoted by $\mathcal{A}^{t}$, is defined as a mapping of resources to agents such that $\mathcal{A}^{t}_{i}$ represents the resources allocated to agent $i$. The welfare corresponding to the post-allocation utility vector $Z^{t}|\mathcal{A}^{t}$ is then given by $W(Z^{t}|\mathcal{A}^{t})$, and the optimal allocation is defined as:
$\mathcal{A}^{t*}$ $=\operatornamewithlimits{argmax}_{\mathcal{A}}W(Z^{t}|\mathcal{A}).$ (5)
@@
There exist various methods to compute the utility vector $Z^{t}|\mathcal{A}$, and these choices influence the resulting allocation. We discuss some popular approaches below, motivating and building up to past-discounted fairness.
@@ 4. Temporal Fairness in Resource Allocation
In dynamic resource allocation, fairness must be evaluated not only on the basis of the current decision but also by considering past allocations and future expectations. In this section, we present three paradigms for temporal fairness: Instantaneous fairness, perfect-recall historical fairness, and discounted-recall historical fairness. We define each approach, illustrate them with examples, and discuss their inherent limitations.
@@ 4.1. Instantaneous Fairness
An allocation exhibits instantaneous fairness if it optimizes a welfare function solely based on the one-step utilities $u_{i}^{t}$.
@@
[…]
In a purely instantaneous allocation, the donut is assigned to the agent who slightly benefits from it more in that step (Alice). Over repeated interactions, however, Alice may receive a disproportionate number of donuts, leading to a cumulative imbalance.
Thus, while instantaneous fairness might maximize short-term efficiency, its disregard for temporal dynamics can result in significant long-term disparities.
@@ 4.2. Perfect-Recall Historical Fairness
To capture the temporal aspect of fairness, perfect-recall historical fairness incorporates all past allocations into the fairness evaluation. Instead of relying solely on the instantaneous utility, we define an adjusted utility vector $Z^{t}$ that aggregates utilities over all previous steps:
$Z^{t}_{i}=\sum_{\tau=0}^{t}u_{i}^{\tau}.$
@@
$Z^{t}_{i}|\mathcal{A}$ $=\sum_{\tau=0}^{t-1}u^{\tau}_{i}+u^{\mathcal{A}}_{i},$
$Z^{t}_{i}|\mathcal{A}$ $=Z^{t-1}_{i}+u^{\mathcal{A}}_{i}.$
@@
$Z^{t}_{i}|\mathcal{A}$ $=\frac{Z^{t-1}_{i}\cdot(t-1)+u^{\mathcal{A}}_{i}}{t}.$
@@
An allocation exhibits perfect-recall fairness if it optimizes a welfare function $W$ over the cumulative (or averaged) utility vector $\bigl{(}Z^{t}_{1},\dots,Z^{t}_{n}\bigr{)}$, where $Z^{t}_{i}$ captures all past allocations.
@@
4.3. Discounted-Recall Historical Fairness
@@ 4.3.1. Discounted Recall with Additive Utilities
$Z^{t}_{i}|\mathcal{A}=\gamma_{p}Z^{t-1}_{i}+u^{\mathcal{A}}_{i}.$ (7)
@@ 4.3.2. Discounted Recall with Averaged Utilities
$Z^{t}_{i}|\mathcal{A}$ $=\frac{\gamma_{p}Z^{t-1}_{i}\cdot d_{t-1}+u^{\mathcal{A}}_{i}}{\gamma_{p}d_{t-1}+1}.$
In both formulations, the computation of $Z^{t}$ depends only on the previous state—specifically, the augmented state comprising $Z^{t-1}$ (and $d_{t-1}$ in the averaged case). This Markovian structure not only simplifies the computation but also provides a smooth interpolation between instantaneous and perfect-recall fairness, while ensuring that the state space is augmented in a tractable manner.
@@ 4.4. Comparison between the Different Paradigms
Figure 1. Comparison of cumulative utility differences under different fairness paradigms. (Left) Cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob}$, over time for Example 2, where both agents participate from the start. (Center) Cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob}$, over time for Example 4, where only Alice is active initially and Bob joins later. (Right) Difference in perceived utility between Alice and Bob for all three methods for Example 4. This plot shows the effect of $\gamma_{p}$ on the perceived values, demonstrating how it changes the speed at which we forget past decisions, interpolating between perfect-recall and instantaneous fairness.
Figure 1 illustrates the evolution of the cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob},$ under the three fairness approaches: Instantaneous, perfect-recall, and discounted-recall fairness, as discussed in Examples 2 and 4. The allocations are made with the MMF objective, using additive aggregation for perfect-recall and discounted-recall. In Figure 1(left), where both agents participate from the start, instantaneous fairness accumulates short-term differences leading to long-term unfairness, while perfect-recall fairness compensates by accounting for all past allocations. Discounted-recall fairness offers a tunable middle ground, with the discount factor $\gamma_{p}$ controlling how quickly past utilities decay in importance.
@@
Finally, Figure 1(right) shows the inner workings of each fairness approach by plotting the perceived differences between Alice and Bob in terms of $Z$. Instantaneous fairness seems to always keep a low difference between the two agents locally, even as the total utility for Alice keeps rising. Perfect-recall keeps an exact track of resources, considering the distribution unfair even after many steps of allocating to Bob only. The decay of the line, particularly visible for $\gamma_{p}=0.9$ shows how discounted-recall slowly forgets past decisions, with values close to 1 emulating longer memory.
4.5. Practical Benefits of Past-Discounting

### v2

# Past-Discounting is Key for Learning Markovian Fairness with Long Horizons

@@ Abstract.
Fairness is an important consideration for dynamic resource allocation in multi-agent systems. Many existing methods treat fairness as a one-shot problem without considering temporal dynamics, which misses the nuances of accumulating inequalities over time. Recent approaches overcome this limitation by tracking allocations over time, assuming perfect recall of all past utilities. While the former neglects long-term equity, the latter introduces a critical challenge: the augmented state space required to track cumulative utilities grows unboundedly with time, hindering the scalability and convergence of learning algorithms. Motivated by behavioral insights that human fairness judgments discount distant events, we introduce a framework for temporal fairness that incorporates past-discounting into the learning problem. This approach offers a principled interpolation between instantaneous and perfect-recall fairness. Our central contribution is a past-discounted framework for memory tracking and a theoretical analysis of fairness memories, showing past-discounting guarantees a bounded, horizon-independent state space, a property that we prove perfect-recall methods lack. This result unlocks the ability to learn fair policies tractably over arbitrarily long horizons. We formalize this framework, demonstrate its necessity with experiments showing that perfect recall fails where past-discounting succeeds, and provide a clear path toward building scalable and equitable resource allocation systems.
@@
Fairness, Resource Allocation, Long-horizon Learning
@@ 1. Introduction
Dynamic resource allocation is central to many real-world applications, from matching passengers to taxis Shah et al. (2020); Alonso-Mora et al. (2017) and distributing aid to the homeless Kube et al. (2019), to allocating life-saving vaccines Alamdari et al. (2024). In these long-horizon settings, agents must balance efficiency with fairness, but existing approaches are caught in a fundamental dilemma. Myopic fairness approaches look at fairness within a single allocation step, ignoring any past allocations and leading to cumulative inequity. To overcome this limitation, recent works tackle temporal fairness by tracking cumulative utilities over time Jiang and Lu (2019); Zimmer et al. (2021); Alamdari et al. (2024)—what we term perfect-recall fairness—but these introduce a critical and previously unaddressed failure mode. As noted by Alamdari et al. (2024), maintaining fairness with history dependence requires augmenting the state. However, we show that with perfect recall, this augmented state space grows unboundedly with time, rendering learning algorithms computationally intractable and preventing them from converging on fair policies.
How can an agent remember the past without being crippled by it? We find inspiration in behavioral economics and moral psychology, which show that human fairness judgments naturally discount distant events. People perceive events in the distant past more abstractly Trope and Liberman (2010), devalue temporally removed outcomes Frederick et al. (2002), and are more inclined to forgive past transgressions as time passes Li et al. (2021); Wohl and McGrath (2007). Motivated by these insights, we introduce a framework that incorporates past-discounting. By applying a tunable discount factor to historical utilities, our approach provides a principled interpolation between flawed instantaneous fairness and intractable perfect-recall fairness. This is not merely a heuristic; we prove that past-discounting is the key to achieving what was previously impossible: Markovian fairness with a provably bounded and horizon-independent state space. This theoretical guarantee ensures the computational tractability required for reinforcement learning agents to learn robustly fair policies, no matter how long the task horizon.
Contributions: Our contributions are:
  1. (1)
    We identify and formalize a critical failure mode in long-term fairness methods Alamdari et al. (2024); Zimmer et al. (2021): perfect-recall causes unbounded state-space growth, rendering learning intractable over long horizons.
  2. (2)
    We propose past-discounted fairness memories, a novel, behaviorally-grounded framework that balances historical context with computational feasibility by applying a geometrically decaying weight to past utilities.
  3. (3)
    We provide a rigorous theoretical proof that our past-discounting mechanism guarantees a bounded and horizon-independent augmented state space, retaining the Markov property while enabling tractable learning.
  4. (4)
    We empirically demonstrate the necessity of discounting the past, showing that an RL agent successfully learns fair policies with our method while systematically failing with perfect recall as the time horizon increases.
@@ 2. Related Work
Fairness in resource allocation has a rich history, but traditional literature has focused on static, one-shot problems. Concepts like proportionality, envy-freeness, and maximin share guarantees Budish (2011); Procaccia and Wang (2014) provide robust solutions for individual decision points. However, their myopic nature makes them insufficient for dynamic settings like taxi matching or aid distribution, where long-term equity necessitates reasoning over multiple time steps. Even in the sequential setting, many approaches consider fairness over only the current valuations of resources or considering future value, but do not consider past memories Sinclair et al. (2022); Hosseini et al. (2024). We call this myopic or instantaneous fairness.
Recognizing this, recent research has explored temporal fairness across various domains. In multi-agent reinforcement learning (MARL), some approaches constrain per-step allocations or evaluate cumulative utilities at a terminal stage Jiang and Lu (2019); Zimmer et al. (2021). More advanced mechanisms include forecasting for ride-hailing applications Kumar et al. (2023) and hierarchical frameworks to mediate between efficiency and fairness Jiang and Lu (2019). Similar challenges are addressed in sequential voting Lackner (2020) and online fair division Aleksandrov et al. (2015), where a series of interdependent choices necessitates a long-term perspective. These works establish a clear consensus: achieving meaningful fairness in sequential settings requires memory of the past.
The dominant paradigm for incorporating this memory is perfect recall, where agents track the complete sum of historical utilities Zimmer et al. (2021); Siddique et al. (2020); Alamdari et al. (2024); Jiang and Lu (2019); Cookson et al. (2025); Lodi et al. (2024). The recent work by Alamdari et al. (2024) formalizes this by treating history-dependent fairness as a non-Markovian problem. They show that by augmenting the state with a memory of past utilities (i.e., perfect recall), one can restore the Markov property. However, this approach introduces a critical, unaddressed challenge: for a perfect memory, the augmented state space grows unboundedly with the time horizon. This makes learning computationally intractable, especially in long-horizon tasks. Our work directly addresses this limitation: we seek a memory mechanism that is not only Markovian but also improves tractability.
This leaves the field caught between two extremes: myopic fairness, which is computationally simple but ignores long-term equity (especially over past allocations), and perfect-recall fairness, which is equitable in theory but can be computationally infeasible, as we show later. Our work introduces a third paradigm that occupies the practical middle ground: past-discounted fairness. This approach is motivated by strong evidence from behavioral economics and moral psychology, which shows that human fairness judgments are sensitive to temporal distance. People naturally devalue outcomes that are further in the past Trope and Liberman (2010); Frederick et al. (2002) and are more forgiving of older transgressions Li et al. (2021); Wohl and McGrath (2007). This concept is analogous to future-discounting in standard reinforcement learning Mnih et al. (2013); Sutton and Barto (1998), but its application to past-oriented fairness memories to ensure tractability remains unexplored.
In summary, while prior work has correctly identified the need for memory in temporal fairness, it has not provided a computationally viable solution for long-horizon problems. Our contribution is to formalize a behaviorally-grounded, past-discounted memory that is (i) sufficient to restore the Markov property with a bounded, horizon-independent state, and (ii) provides a practical mechanism for learning fair policies in complex, long-running systems. To the best of our knowledge, no existing work has considered using past discounts to learn fair behavior.
@@ 3. Preliminaries
    $W_{N}(\mathbf{Z})=\prod_{i=1}^{n}z_{i},$ (4)
@@
In many formulations, the allocation at time $t$, denoted by $\mathcal{A}^{t}$, is defined as a mapping of resources to agents such that $\mathcal{A}^{t}_{i}$ represents the resources allocated to agent $i$. The welfare corresponding to the post-allocation utility vector $\mathbf{Z}^{t}|\mathcal{A}^{t}$ is then given by $W(\mathbf{Z}^{t}|\mathcal{A}^{t})$, and the optimal allocation is defined as:
$\mathcal{A}^{t*}$ $=\operatornamewithlimits{argmax}_{\mathcal{A}}W(\mathbf{Z}^{t}|\mathcal{A}).$ (5)
@@
There exist various methods to compute the utility vector $\mathbf{Z}^{t}|\mathcal{A}$, and these choices influence the resulting allocation. We discuss some popular approaches below, motivating and building up to past-discounted fairness.
Figure 1. Comparison of cumulative utility differences under different fairness paradigms. (Left) Cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob}$, over time for Example 2, where both agents participate from the start. (Center) Cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob}$, over time for Example 4, where only Alice is active initially and Bob joins later. (Right) Difference in perceived utility between Alice and Bob for all three methods for Example 4. This plot shows the effect of $\gamma_{p}$ on the perceived values. It changes the speed at which we forget past decisions, interpolating between perfect-recall and instantaneous fairness.
@@ 4. Temporal Fairness in Resource Allocation
In dynamic resource allocation, fairness must be evaluated not only on the basis of the current decision but also by considering past allocations and future expectations. In this section, we present three paradigms for temporal fairness: Instantaneous fairness, perfect-recall historical fairness, and our proposed discounted-recall historical fairness. We define each approach, illustrate them with examples, and discuss their inherent limitations.
@@ 4.1. Instantaneous Fairness
An allocation exhibits instantaneous fairness at time $t$ if fairness is evaluated solely on the one-step utilities. For any candidate allocation $\mathcal{A}$, define
$\mathbf{Z}^{t}\mid\mathcal{A}=\bigl(u^{\mathcal{A}}_{1},\ldots,u^{\mathcal{A}}_{n}\bigr).$ (7)
Then an instantaneous-fair allocation at time $t$ is any optimizer that selects the allocation that maximizes the welfare applied to the current-step utility vector.
@@
$\text{Alice: }(u_{\text{cake}},u_{\text{donut}})=(0.2,0.5),\quad\text{Bob: }(0.3,0.5).$
In a purely instantaneous allocation, the donut is assigned to the agent who slightly benefits from it more in that step (Alice, compared to the baseline 0.2). Over repeated interactions, however, Alice may receive a disproportionate number of donuts, leading to a cumulative imbalance.
Thus, while instantaneous fairness might maximize short-term objectives, its disregard for temporal dynamics can result in significant long-term disparities.
@@ 4.2. Perfect-Recall Historical Fairness
To capture the temporal aspect of fairness, perfect-recall historical fairness incorporates all past allocations into the fairness evaluation. Instead of relying solely on the instantaneous utility, we define an adjusted utility vector $\mathbf{Z}^{t}$ that aggregates utilities over all previous steps:
$Z^{t}_{i}=\sum_{\tau=0}^{t}u_{i}^{\tau}.$ (8)
@@
$Z^{t}_{i}\mid\mathcal{A}$ $=\sum_{\tau=0}^{t-1}u^{\tau}_{i}+u^{\mathcal{A}}_{i}$ (9)
$=Z^{t-1}_{i}+u^{\mathcal{A}}_{i}.$ (10)
@@
$Z^{t}_{i}\mid\mathcal{A}=\frac{Z^{t-1}_{i}\cdot(t-1)+u^{\mathcal{A}}_{i}}{t}.$ (11)
@@
An allocation exhibits perfect-recall fairness at time $t$ if fairness is evaluated using all past allocations. For any candidate allocation $\mathcal{A}$, define $\mathbf{Z}^{t}\mid\mathcal{A}$ by either:
(cumulative) $Z^{t}_{i}\mid\mathcal{A}=Z^{t-1}_{i}+u^{\mathcal{A}}_{i},$ (12)
(averaged) $Z^{t}_{i}\mid\mathcal{A}=\frac{Z^{t-1}_{i}\cdot(t-1)+u^{\mathcal{A}}_{i}}{t}.$ (13)
A perfect-recall-fair allocation at time $t$ is any optimizer that selects the allocation that maximizes the welfare applied to the history-aggregated utility vector.
@@
4.3. Past-Discount: Discounted-Recall Historical Fairness
@@ 4.3.1. Discounted Recall with Additive Utilities
$Z^{t}_{i}\mid\mathcal{A}=\gamma_{p}\,Z^{t-1}_{i}+u^{\mathcal{A}}_{i}.$ (14)
@@
Definition 0 (Discounted-Recall Fairness (Additive)).
An allocation exhibits discounted-recall fairness (additive) at time $t$ if fairness is evaluated using the past-discounted update. For any candidate allocation $\mathcal{A}$, define
$Z^{t}_{i}\mid\mathcal{A}=\gamma_{p}\,Z^{t-1}_{i}+u^{\mathcal{A}}_{i}\quad\text{for }\gamma_{p}\in[0,1].$ (15)
A discounted-recall-fair (additive) allocation at time $t$ is any optimizer that selects the allocation that maximizes the welfare applied to the past-discounted cumulative utilities.
@@ 4.3.2. Discounted Recall with Averaged Utilities
$Z^{t}_{i}\mid\mathcal{A}=\frac{\gamma_{p}Z^{t-1}_{i}\cdot d_{t-1}+u^{\mathcal{A}}_{i}}{\gamma_{p}d_{t-1}+1}.$ (16)
Definition 0 (Discounted-Recall Fairness (Averaged)).
An allocation exhibits discounted-recall fairness (averaged) at time $t$ if fairness is evaluated using the past-discounted average with denominator $d_{t}$. For any candidate allocation $\mathcal{A}$, define
$Z^{t}_{i}\mid\mathcal{A}=\frac{\gamma_{p}Z^{t-1}_{i}\cdot d_{t-1}+u^{\mathcal{A}}_{i}}{\gamma_{p}d_{t-1}+1}.$ (17)
A discounted-recall-fair (averaged) allocation at time $t$ is any optimizer that selects the allocation that maximizes the welfare applied to the past-discounted average utilities.
In both formulations, the computation of $\mathbf{Z}^{t}$ depends only on the previous state—specifically, the augmented state comprising $\mathbf{Z}^{t-1}$ (and $d_{t-1}$ in the averaged case). This Markovian structure not only simplifies the computation but also provides a smooth interpolation between instantaneous and perfect-recall fairness, while ensuring that the state space is augmented in a tractable manner.
@@ 4.4. Comparison between the Different Paradigms
Figure 1 illustrates the evolution of the cumulative utility difference, $\sum U_{Alice}-\sum U_{Bob},$ under the three fairness approaches: Instantaneous, perfect-recall, and discounted-recall fairness, as discussed in Examples 2 and 4. The allocations are made with the egalitarian objective, using additive aggregation for perfect-recall and discounted-recall. In Figure 1(left), where both agents participate from the start, instantaneous fairness accumulates short-term differences leading to long-term unfairness, while perfect-recall fairness compensates by accounting for all past allocations. Discounted-recall fairness offers a tunable middle ground, with the discount factor $\gamma_{p}$ controlling how quickly past utilities decay in importance.
@@
Finally, Figure 1(right) shows the inner workings of each fairness approach by plotting the perceived differences between Alice and Bob in terms of $\mathbf{Z}$. Instantaneous fairness seems to always keep a low difference between the two agents locally, even as the total utility for Alice keeps rising. Perfect-recall keeps an exact track of resources, considering the distribution unfair even after many steps of allocating to Bob only. The decay of the line, particularly visible for $\gamma_{p}=0.9$ shows how discounted-recall slowly forgets past decisions, with values close to 1 emulating longer memory.
5. Theoretical Results
So far, we have established qualitative and behavioral motivations for the three approaches through illustrative examples. In this section, we present theoretical results that highlight the practical advantages of past-discounted fairness over perfect-recall fairness. We first focus on the boundedness of the augmented state space, which is crucial for the convergence and scalability of reinforcement learning algorithms, especially in long-horizon settings.
5.1. Additive Utilities